# Thermodynamics of black and white holes in ensemble of Planckons

### v4

@@ III.1 Probability of formation of white hole and negative entropy
This equation suggests the analog of Bekenstein entropy boundBekenstein1981 ; Bekenstein2004 and holographic entropy boundHooft1993 ; Susskind1995 , which now includes the negative entropy of white holes. The entropy of an ensemble of $N$ Planckons is bounded by maximum and minimum values that correspond to the entropies of black and white holes:
@@
The entropy of a white hole is the negative entropy of the least probable exceptional state. A white hole relaxes into a black hole with increasing entropy, and thus due to the negative entropy of a white hole, the second law of thermodynamics is not violated.
Eq. (18) is supported by the direct calculations of the transition from the black hole to the white hole by macroscopic quantum tunneling,Volovik2022 which demonstrate that the entropy of the white hole with mass $M$ is with minus sign the entropy of the black hole with the same mass:
$S_{\rm WH}(M)=-S_{\rm BH}(M)=-4\pi GM^{2}\,.$ (20)
In the same manner the time reversal transforms the expanding de Sitter with the positive Hubble parameter, $H>0$, to contracting de Sitter with $H<0$. Correspondingly, due to anti-symmetry of the Hubble parameter, $TH=-H$, the entropy of the volume inside the cosmological horizon (the Hubble volume) is negative in the contracting Universe, $S(H<0)=-S(H>0)=-A/4G$, where $A$ is the area of the cosmological horizon.Volovik2025b In transition between expanding and contracting Universe via Minkowskii state, the local entropy density $s$ changes sign by crossing zero value, which is natural for the empty Minkowski vacuum. But the total entropy $S(H)=sV_{H}$ of the Hubble volume $V_{H}$ changes sign by crossing an infinite value. This shows that spatial infinity in the Minkowski vacuum can play a role similar to that of the cosmological event horizon. Escape to infinity may play the same role as escape through the crossing of the cosmological horizon.
III.2 Tsallis-Cirto statistics with negative entropy and second law of thermodynamics
@@ IV.3 Horizons, thermodynamics and Lifshitz transition
The equation (29) is also applicable to the processes associated with the cosmological horizon. The corresponding temperature governing the processes of radiation from the horizon is the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$, where $H$ is the Hubble parameter. The local temperature of the de Sitter state, describing the de Sitter thermodynamics, is not related to horizon and is twice as large.Volovik2025b Nevertheless, the local entropy density $s$ integrated over the Hubble volume $V_{H}$ (the volume inside the cosmological horizon) corresponds to the Gibbons-Hawking entropy of the cosmological horizon, $S_{H}=sV_{H}=A/4G$, where $A$ is the area of the cosmological horizon. This connection between local thermodynamics and the thermodynamics of the cosmological horizon is a consequence of the specific symmetry of the de Sitter state and is supported by topological arguments.Kusmartsev2025 ; Volovik2025f
Another connection is provided by the Planckon model. According to this model, the energy and entropy of the Hubble volume are quantized in the same way as in the case of a black hole:
@@
$S_{H}(N\gg 1)\rightarrow\frac{1}{2}N^{2}=\frac{\pi}{GH^{2}}=\frac{A}{4G}\,,$ (31)
On the other hand, quantized entropy implies quantization of the cosmological constant $\Lambda$, i.e. the energy density $\rho_{\rm vac}$ of the de Sitter state, which is a local quantity:
@@
Here $l_{\rm P}=1/m_{\rm P}$ is the reduced Planck length. A similar quantization of $\Lambda$ is obtained in the gravitational-topological sector.Wieland2011 ; Alexander2025 Equation (23) in Ref.Wieland2011 implies for the reduced cosmological constant the values $\Lambda=\frac{12\pi}{\beta l_{\rm P}^{2}n}$, where $\beta$ is the Barbero–Immirzi parameter and $n$ is an integer. With $\beta=\frac{1}{2\pi}$ this agrees with Eq.(32).
@@ VI Conclusion
Thermodynamics of black and white holes can be considered in terms of the ensemble of the Planck mass objects, Planckons. The Tsallis-Cirto non-extensive statistics with $\delta=2$, which describes the processes of splitting and merging of black holes, suggests the statistical properties of this ensemble. The entropy of black hole with $N$ Planckons is provided by the $N(N-1)/2$ degrees of freedom – the correlations between the gravitationally attracted Planckons. If one chooses the mass of Planckon as the reduced Planck mass, $m_{\rm P}=1/\sqrt{8\pi G}$, and thus the mass of the black hole is $M=Nm_{\rm P}$, then the entropy of this black hole is $S_{\rm BH}(N)=N(N-1)/2$.
@@
According to Unruh,Unruh2024 the quantum mechanics of black holes is still a far from finished topic. Also according to Jacobson,Jacobson2024 the connection between the de Sitter horizon entropy and quantum field theory still stands on shaky ground. This toy model will likely give us a deeper understanding of the origins of thermodynamics that arise in black holes, as well as in the quantum vacuum.

### v5

@@ II.2 Entropy in terms of Planck-scale elements
The zero entropy of an individual Planckon distinguishes these toy objects from physical Planck-scale black holes.
@@ III.1 Probability of formation of white hole and negative entropy
Eq. (18) is supported by the direct calculations of the transition rate from the black hole to the white hole by macroscopic quantum tunneling,Volovik2022 which demonstrate that the entropy of the white hole with mass $M$ is with minus sign the entropy of the black hole with the same mass:
$S_{\rm WH}(M)=-S_{\rm BH}(M)=-4\pi GM^{2}\,.$ (19)
III.2 White hole and modified Bekenstein entropy bound
Eq. (18) suggests the analog of Bekenstein entropy boundBekenstein1981 ; Bekenstein2004 and holographic entropy boundHooft1993 ; Susskind1995 , which now includes the negative entropy of white holes. The entropy of an ensemble of $N$ Planckons is bounded by maximum and minimum values that correspond to the entropies of black and white holes:
@@ III.2 White hole and modified Bekenstein entropy bound
The entropy of a white hole is the negative entropy of the least probable exceptional state. The white hole is unstable to the radiation of Planckons, and in the process of evolution all Planckons leave the white hole. Then these Planckons are attracted to each other and eventually form a black hole. We assume here that the system is finite, so that Planckons cannot escape to infinity after Hawking radiation, otherwise the black hole will finally decay.
In this scenario, as in other scenarios of evolution, the white hole relaxes into a black hole with increasing entropy. This demonstrates that due to the negative entropy of the white hole, the second law of thermodynamics is not violated. On the other hand, the inverse process, i.e. the process of formation of white hole, corresponds to quantum fluctuation with probability $w\sim\exp\left(-N(N-1)\right)$.
III.3 Tsallis-Cirto statistics with negative entropy and second law of thermodynamics
@@ IV.3 Horizons, thermodynamics and Lifshitz transition
V.1 Thermodynamics of cosmological horizon
In the Painleve-Gullstrand coordinates, the corresponding metric has the shift vector:
${\bf v}=H{\bf r}\,,$ (30)
where $H$ is the Hubble parameter.
The processes associated with the cosmological horizon at $R=1/H$ also obey the balance equation (29). The corresponding temperature governing the processes of radiation from the horizon is the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$. However, the real local temperature of the de Sitter state which describes the de Sitter thermodynamics, is not related to horizon and is twice as large, $T=H/\pi=2T_{\rm GH}$.Volovik2025b This local temperature determines the local entropy density $s$ of the de Sitter state. The total entropy of the Hubble volume $V_{H}$ (the volume inside the cosmological horizon) nevertheless coincides with the Gibbons-Hawking entropy $S_{\rm GH}$ of the cosmological horizon:
$S_{H}=sV_{H}=\frac{A}{4G}=S_{\rm GH}\,,$ (31)
where $A$ is the area of the cosmological horizon. Thus, although the local temperature does not coincides with the Gibbons-Hawking temperature, this does not violate the holographic bulk-surface correspondence. The factor two between the local temperature $T$ and the temperature $T_{\rm GH}$ associated with the cosmological horizon is a consequence of a special symmetry of the de Sitter state and is supported by topological arguments.Kusmartsev2025 ; Volovik2025f
Another connection is provided by the Planckon toy model. According to this model, the energy and entropy of the Hubble volume are quantized in the same way as in the case of a black hole:
@@ V.1 Thermodynamics of cosmological horizon
$S_{H}(N\gg 1)\rightarrow\frac{1}{2}N^{2}=\frac{\pi}{GH^{2}}=\frac{A}{4G}\,.$ (33)
V.2 Quantization of cosmological constant
While the quantized entropy of the Hubble volume is a rather abstract quantity, the Planckon model suggests also the quantization of the physical observable – the cosmological constant $\Lambda$. This is the energy density $\rho_{\rm vac}$ of the de Sitter state, which is a local quantity. From Planckon model one obtains:
@@ V.2 Quantization of cosmological constant
Here $l_{\rm P}=1/m_{\rm P}$ is the reduced Planck length. A similar quantization of $\Lambda$ is obtained in the gravitational-topological sector.Wieland2011 ; Alexander2025 Equation (23) in Ref.Wieland2011 implies for the reduced cosmological constant the values $\Lambda=\frac{12\pi}{\beta l_{\rm P}^{2}n}$, where $\beta$ is the Barbero–Immirzi parameter and $n$ is an integer. With $\beta=\frac{1}{2\pi}$ and $n=\cal N$, this agrees with Eq.(34).
@@
V.3 Negative entropy of contracting de Sitter
The time reversal transforms the expanding de Sitter with the positive Hubble parameter, $H>0$, to contracting de Sitter with $H<0$. Correspondingly, due to anti-symmetry of the Hubble parameter, $TH=-H$, the entropy of the volume inside the cosmological horizon (the Hubble volume) is negative in the contracting Universe,Volovik2025b
$S(H<0)=-S(H>0)=-\frac{A}{4G}\,.$ (35)
Transition between expanding and contracting states of de Sitter Universe may occur via Minkowski state. In this scenario, the local entropy density $s$ changes sign by crossing the zero value, which is natural for the empty Minkowski vacuum with temperature $T=0$. But at the transition point, i.e. in Minkowski vacuum, the horizon is at $r=\infty$. This means that in this transition, the total entropy $S(H)$ of the Hubble volume $V_{H}$ changes sign by crossing an infinite value in the Minkowski vacuum. This demonstrates that spatial infinity in the Minkowski vacuum can play a role similar to that of the cosmological event horizon. Escape to infinity looks similar to escape through the crossing of the cosmological horizon.
The Planckon model also suggests that Minkowski vacuum is not empty. While the density of Planckons $N/V_{H}=3H^{2}m_{\rm P}$ approaches zero in the Minkowski limit, their total number in the Hubble volume, $N=\frac{4\pi m_{\rm P}}{H}$, approaches infinity at $H\rightarrow 0$ (see also Ref.Carr2025 on the decay of de Sitter space into real physical Planckons). This is the consequence of the holographic bulk-surface correspondence. The holography and possible escape to space infinity raise the problem of decoherence in the Minkowski vacuum,Danielson2023 ; Danielson2025 and probably are important for the problem of the collapse of the wave function.Bassi2013 ; Bose2025
@@ VI Conclusion
Thermodynamics of black and white holes can be considered in the toy model of the ensemble of the Planck mass objects, Planckons. The Tsallis-Cirto non-extensive statistics with $\delta=2$, which describes the processes of splitting and merging of black holes, suggests the statistical properties of this ensemble. The entropy of black hole with $N$ Planckons is provided by the $N(N-1)/2$ degrees of freedom – the correlations between the gravitationally attracted Planckons. If one chooses the mass of Planckon as the reduced Planck mass, $m_{\rm P}=1/\sqrt{8\pi G}$, and thus the mass of the black hole is $M=Nm_{\rm P}$, then the entropy of this black hole is $S_{\rm BH}(N)=N(N-1)/2$.
@@
According to Unruh,Unruh2024 the quantum mechanics of black holes is still a far from finished topic. Also according to Jacobson,Jacobson2024 the connection between the de Sitter horizon entropy and quantum field theory still stands on shaky ground. The Planckon toy model will likely give us a deeper understanding of the origins of thermodynamics that arise in black holes, as well as in the quantum vacuum.